# CGRA4ML: A Framework to Implement Modern Neural Networks for Scientific Edge Computing

### v2

@@
Abstract
Scientific edge computing increasingly relies on hardware-accelerated neural networks to implement complex, near-sensor processing at extremely high throughputs and low latencies. Existing frameworks like hls4ml are effective for smaller models, but struggle with larger, modern neural networks due to their requirement of spatially implementing the neural network layers and storing all weights in on-chip memory. CGRA4ML is an open-source, modular framework designed to bridge the gap between neural network model complexity and extreme performance requirements. CGRA4ML extends the capabilities of hls4ml by allowing off-chip data storage and supporting a broader range of neural network architectures, including models like ResNet, PointNet, and transformers. Unlike hls4ml, CGRA4ML generates SystemVerilog RTL, making it more suitable for targeting ASIC and FPGA design flows. We demonstrate the effectiveness of our framework by implementing and scaling larger models that were previously unattainable with hls4ml, showcasing its adaptability and efficiency in handling complex computations. CGRA4ML also introduces an extensive verification framework, with a generated runtime firmware that enables its integration into different SoC platforms. CGRA4ML’s minimal and modular infrastructure of Python API, SystemVerilog hardware, Tcl toolflows, and C runtime, facilitates easy integration and experimentation, allowing scientists to focus on innovation rather than the intricacies of hardware design and optimization.
I Introduction
Figure 1: CGRA4ML workflow as outlined in Sec. III. Users first build quantized neural networks and train them in a quantization-aware manner using qkeras (Sec. V-B). Users then define a CGRA definition, which generates vendor-agnostic SystemVerilog RTL hardware specifications and TCL tool flows (Sec. IV). The model can then be exported to generate weights and a C runtime firmware (Sec. V-C). The generated hardware IP is then verified comprehensively with the model and firmware (Sec. V-D) using our randomized, transactional SystemVerilog testbench suite with DPI-C extensions. Finally, the bitstream generated from the FPGA toolflow and the C firmware can be loaded into an FPGA to be tested in seconds (Sec. V-E1). After such rapid prototyping, the same hardware design can be moved to ASIC (Sec. V-E2).
Scientific applications have unique and challenging constraints to process the data quickly and closely to where the data is generated [1]. This enforces strict real-time requirements to process the information at tremendously high throughputs (TB/s) and low latencies. Thus, fast machine-learning scientific applications often require hardware accelerators implemented on FPGAs or ASICs to meet the stringent performance requirements [2].
hls4ml is a popular tool for developing hardware-accelerated scientific neural networks, which implements a dataflow-style architecture with layer-specific datapaths and on-chip weight storage [3]. While this methodology provides high throughput and low latency, it does not support modern, deep neural network models (DNNs) that are too big to fit fully on-chip, requiring off-chip movement of weights and data. In addition, modern neural network models require the development of optimized layer architectures, e.g., a transformer cannot easily be implemented in hls4ml due to its intensive customization and large memory requirements for weights. Most accelerators presented in the literature that support DNNs do not offer usable frameworks with parametrization of their architecture or provide support for verification and SoC integration, as described in Sec. V. The frameworks available for accelerating larger models in FPGAs do not fulfill the requirements of the scientific computing community, such as more robust quantization, extensibility to add more features, and the need to eventually move from FPGA to ASIC implementation, as summarized in Table I.
CGRA4ML bridges the gap between hls4ml and the needs of the scientific community to implement modern, large neural network models using an easy-to-use, high-level workflow. Like hls4ml, CGRA4ML is entirely open source, allowing the scientific community to use, develop, and build upon it. CGRA4ML is easily programmable and scalable, allowing seamless scaling to fit the resources at hand while being easy and intuitive to integrate into the fast design prototyping flow. CGRA4ML generates a parameterizable Coarse-Grained Reconfigurable Array (CGRA) to facilitate the spatial reuse of processing elements and leverage off-chip data storage to deploy modern neural network features. Our Python API works on top of qkeras, a popular library for quantizing and training neural networks. CGRA4ML emits RTL instead of HLS code, making it more amenable to ASIC and FPGA design flows compared to hls4ml.
CGRA4ML builds and configures parametrized flexible CGRA that is then programmed to compute different neural network models. CGRA4ML facilitates resource sharing across neural network layers, as opposed to hls4ml, which implements each layer as a separate datapath. CGRA4ML scales beyond hls4ml capabilities, e.g., implementing larger ResNet models that fail to synthesize with hls4ml. This enables scientific applications that require more complex models that cannot fit within an FPGA or given area of an ASIC. We demonstrate how CGRA4ML can be extended to handle different neural network architectures, including ResNets, PointNet, and transformers. CGRA4ML implements its hardware as fully parametrized SystemVerilog RTL modules; the same implementation is supported in different FPGA and ASIC tools.
@@ I Introduction
CGRA4ML is targeted to two user groups: (1) Scientists, embedded / IoT researchers who want to build and implement modern neural networks that are infeasible in hls4ml due to their size and complexity. (2) Hardware system designers who want to iteratively test new optimizations such as new datatypes, dataflow, and compression techniques.
CGRA4ML provides a Python API for the end-to-end workflow to allow scientists to build and implement their designs to production-ready quality without spending months on designing and verifying hardware, optimizing computation and data bottlenecks, integrating into an SoC, and writing, verifying, and optimizing runtime firmware (see Section III). The CGRA4ML infrastructure, shown in Fig 1 and described in Section V, is modular and extensible to allow the hardware researchers to modify any part, implement their optimizations, and get it working without building the entire system and optimizing every part of it themselves.
To demonstrate the capabilities of CGRA4ML and to compare with other tools our scientific users are familiar with, we present the implementations of ResNet-50, PointNet, and Particle Transformer models in Section VI. We also present the performance, power, and area comparisons of the FPGA and ASIC implementations generated by CGRA4ML.
The major contributions of CGRA4ML are:
@@
  • A CGRA architecture with dynamic reconfiguration, unified dataflow, and an efficient PE design to maximize hardware utilization for modern neural networks.
@@
  • Automated firmware generation enabling hardware/ software partitioning that easily integrates the CGRA with on-chip processors. Our results show how to partition DNN layers across an ARM core and the CGRA.
  • Validation of results on AMD Xilinx FPGAs and ASIC implementation using Cadence tools to be taped out as an SoC with ARM-based NanoSoC platform.
@@
II Background and Related Work
II-A Neural Networks for Scientific Computing
@@
Machine learning is fueling scientific discoveries in particle physics, materials, dark matter, cosmology, nuclear physics, biomedical engineering, and health monitoring [2]. Specific examples include bio-signal classification [4], tracking magnetohydrodynamic (MHD) instabilities in fusion reactors [5], reinforcement learning for accelerator beam control [6], pulse detection for anti-neutrino detection [7], and accelerator controls for beam loss deblending [8].
Scientific computing increasingly relies on neural networks to process data with requirements for extremely low latencies and high throughputs. For example, the Large Hadron Collider (LHC) has $\mathcal{O}(100M)$ individual sensors generating data for each of the 40 million proton beam collisions per second [9]. Neural network-based autoencoders implemented as ASICs compress data to move it off chip [10]. LHC trigger systems filter events in real-time with sub-microsecond latency requirements and 40 MHz throughput requirements using FPGAs [11]
The scientific community has benefited greatly from the ease of hardware acceleration. Among the multiple tools [12] that allow FPGA implementation of neural networks, hls4ml is very popular [13]. Yet, hls4ml has drawbacks as discussed in the next section. CGRA4ML allows more complex models to be implemented on FPGA and ASIC-based SoCs, moving more computation to the edge. This enables more sophisticated neural networks to be deployed in scientific equipment.
II-B hls4ml
hls4ml is a Python library developed by physicists for machine learning inference in FPGAs [13]. It primarily supports models built and trained using QKeras, a library from Google for low-precision machine learning [14]. hls4ml converts the model, layer by layer into customized high-level synthesis (HLS) hardware blocks, as shown in Figure 2. hls4ml uses different HLS implementations for different HLS tools to generate the hardware accelerator depending on the resulting target (AMD Xilinx FPGA, Altera FPGA, ASIC, etc.). The most common target is AMD Xilinx Vitis/Vivado HLS, though other backends exist in various forms (Intel HLS Compiler, Siemens Catapult HLS, etc.). The dependence on HLS restricts the user to the supported vendors and requires HLS tool-specific reimplementation of each hardware layer, complicating the maintenance and verification. hls4ml’s support for ASIC implementation is primarily through Catapult HLS, which currently only supports a limited subset of layers.
The most common hls4ml design flow is to generate an overall IP for the entire neural network and then integrate that IP into the SoC using vendor-specific IPs such as DMAs. There are efforts for better end-to-end development. For example, hls4ml provides a PYNQ overlay, a Python-based runtime for AMD Xilinx Zynq FPGAs. This workflow allows the physicists to focus on building and training low-precision models for their custom applications without spending months on hardware design, implementation, and verification. Yet, by and large, the ML IP core integration for production-ready designs is a separate process from hls4ml.
hls4ml implements the neural network layer by layer, as shown in Fig. 2. The user controls reuse factor per layer, which defines the number of times a multiplier/accumulator is reused. The resources needed to synthesize a layer and the number of clock cycles before its hardware block can accept new data both depend on the reuse factor. Commonly, the same reuse factor is used for all layers. This results in heavy underutilization since all hardware blocks must wait for the slowest to complete. Manual tuning of the reuse factors per layer on each target FPGA to balance the dataflow takes a lot of time and effort.
While the layer-by-layer approach can offer in low-latency, efficient implementation of small models that fully fit into FPGAs, it does not support modern neural network models like transformers. There have been herculean efforts to implement models like ResNet-50 on up to 12 interconnected FPGAs [15]. Other large models, e.g., jet tagging using a transformer [9], are implemented on GPUs because they are impossible to implement with hls4ml. The data filtering pipeline from the detectors of the Large Hadron Collider at CERN implements a tiny autoencoder NN on an ASIC closest to the detector. The data from the ASIC is then processed by progressively more complex NNs on FPGAs and servers. CGRA4ML strives to provide a user-friendly, more customizable workflow similar to hls4ml while supporting larger DNNs, allowing more computations to be moved to the edge and improving the accuracy of the data filtering pipeline.
@@
Figure 2: Layer-by-layer spatial implementation of neural networks using hls4ml. hls4ml offers multiple backends, one per family of devices. Each backend is a collection of layers defined as templated HLS. Each neural network layer is implemented as a separate datapath. Xilinx DMA IPs are connected to the HLS4ML design to move data in and out.
II-C CGRA
CGRAs were originally envisioned as coarse-grained FPGAs whose programmable elements work at the word-level instead of FPGA bit-level programming [16]. Early CGRAs were classified based on their integration into the processor. Tightly coupled CGRAs integrate into a processor data path and are executed as a custom instruction, e.g., Chess [17], MATRIX [18], and DySer [19]. Loosely coupled CGRAs act more as an accelerator that executes alongside the processor, executing in tandem with the processor and communicating via on-chip interconnect. Examples of loosely-coupled CGRAs include PipeRench [20], MorphoSys [21], CHARM [22], and FPCA [23]. CGRA4ML can be characterized as a loosely coupled CGRA as it implements sub-tasks of the neural network alongside a CPU with a focus on enabling hardware/software partitioning across the processor and processor.
@@ II-C CGRA
CGRAs have seen a recent resurgence in industry and academic projects. Commercial CGRAs include Samsung Reconfigurable Processor [24] and the Renesas STP Reconfigurable Processor [25]. CGRA-ME [26] is an exemplary academic framework that supports a subset of CGRA designs.
CGRA4ML is targeted to optimize the neural networks for scientific applications with high throughput, low latency, and real-time requirements. As a result, the default CGRA design of CGRA4ML features small processing elements (PEs) to minimize resource use and an on-the-fly reconfiguration mechanism to reduce the reconfiguration overhead.
This hls4ml VTA FINN Vitis AI DLP OpenVino Caffeine CGRA-ME FlexCNN
work [13] [27] [28] [29] [30] [31] [32] [26] [33]
@@
TABLE I: Comparing CGRA4ML with existing ML to FPGA/ASIC frameworks.
II-D DNN Reconfigurable Architectures
Hundreds of accelerator architectures have been designed to accelerate deep neural networks in the past decade [34, 35, 36]. These accelerators can be categorized by their data reuse pattern as weight, input, output & row stationary and by their runtime flexibility into systolic arrays, CGRAs, and microcode processors.
We compare the CGRA4ML with state-of-the-art architectures. Caffeine [32] offers an HLS systolic array to process CNNs from the Caffe ML framework using weight major and input major mappings. Eyeriss [37] is an energy-efficient ASIC designed to accelerate CNNs using a 2D array of 168 fairly complex processing elements, each with a 16-bit multiply-accumulate, scratchpad, and a controller. ShiDianNao [38] uses a 2D mesh of functional units optimized towards the 2D feature maps of convolutional layers. Each processing element executes multiplications, additions, and comparisons using 16-bit fixed point arithmetic. The kernel elements are shifted right to left and up to down and accumulate locally.
While systolic arrays with simple PEs are optimal for matrix multiplications, they are underutilized and incur a significant overhead when computing more complex layers. CGRAs that allow greater flexibility in data routing optimally utilize their PEs but are able to fit fewer PEs within given resource constraints due to the complexity of the PEs, local scratchpad memories, and routing logic. CGRA4ML aims to strike a balance between these by providing a default CGRA engine with an efficient PE design to maximize PE utilization while allowing flexibility to dynamically regroup to process layers such as convolution maximally exploiting data reuse on multiple levels (see Section IV). CGRA4ML is primarily output stationary; weights and input pixels are stored in on-chip memories to maximize utilization.
II-E ML to FPGA/ASIC Frameworks

### v3

# cgra4ml: A Hardware/Software Framework to Implement Neural Networks for Scientific Edge Computing

@@
Abstract.
The scientific community increasingly relies on machine learning (ML) for near-sensor processing, leveraging its strengths in tasks such as pattern recognition, anomaly detection, and real-time decision-making. These deployments demand accelerators that combine extremely high performance with programmability, ease of integration, and straightforward verification. We present cgra4ml, an open-source, modular framework that generates parameterizable CGRA accelerators in synthesizable SystemVerilog RTL, tailored to common ML compute patterns found in scientific applications. The framework supports seamless system integration through AXI-compliant interfaces and open-source DMA components, and it includes automatic firmware generation for programming the accelerator. A comprehensive verification suite and a runtime firmware stack further support deployment across diverse SoC platforms. cgra4ml provides a modular, full-stack infrastructure, including a Python API, SystemVerilog hardware, TCL toolflows, and a C runtime, which facilitates easy integration and experimentation, allowing scientists to focus on innovation rather than dealing with the intricacies of hardware design and optimization. We demonstrate the effectiveness of cgra4ml to implement common scientific edge neural networks using ASIC and FPGA design flows.
FPGA, reconfigurable computing, edge computing, machine learning, CGRA
journal: TRETSccs: Hardware Reconfigurable logic applicationsccs: Computing methodologies Machine learning
1. Introduction
Modern scientific discovery increasingly relies on machine learning to process vast volumes of data at the edge, where sensors capture events in real-time and decisions must be made within microseconds (Duarte et al., 2018). From high-energy physics experiments and remote sensing arrays to in-situ biomedical diagnostics and fusion plasma control, these scientific applications demand neural network inference on custom hardware architectures to achieve the required inference speed and energy efficiency with extremely high throughput, low latencies, and tight energy budgets (Coelho et al., 2021). Such custom hardware designs must be closely integrated with host microprocessors and memory subsystems to achieve the highest performance. At the same time, these workloads increasingly utilize complex, evolving neural network models, necessitating a degree of programmability and architectural flexibility. Previous scientific edge computing frameworks such as hls4ml (FastML Team, 2024) and finn (Umuroglu et al., 2017) have enabled small neural networks on FPGAs through high-level synthesis. However, they struggle to scale to deeper models, offer limited flexibility in hardware reuse, and lack production-ready verification and integration flows. Therefore, the scientific computing community is in need of a framework that supports larger models with a user-friendly interface, support for various quantization levels, modular hardware architecture, production-ready firmware generation, complete system integration, and verification.
To address these requirements, we present cgra4ml, an open-source framework for deploying modern neural networks in scientific edge computing environments. cgra4ml generates parametrizable, reusable Coarse-Grained Reconfigurable Array (CGRA) hardware engines and associated runtime systems, enabling scalable inference of complex models across FPGAs and ASICs. The framework delivers a holistic solution, comprising a Python-based frontend for quantized model specification, a SystemVerilog RTL backend for the CGRA, a modular C runtime firmware, and an integrated verification suite.
@@ 1. Introduction
cgra4ml is targeted to two user groups: (1) Scientists, embedded / IoT researchers who want to build and implement modern neural networks that cannot be deployed using current popular frameworks for scientific computing due to their size and complexity. (2) Hardware system designers who want to iteratively test new optimizations such as new datatypes, dataflow, and compression techniques.
The major contributions of cgra4ml include:
@@
  • An open-source CGRA architecture with dynamic reconfiguration, unified dataflow, and an efficient PE design to maximize hardware utilization for modern neural networks.
@@
  • Automated firmware generation enabling hardware/software partitioning that easily integrates the CGRA with on-chip processors.
  • End-to-end implementation and test on AMD Xilinx FPGAs and ASIC flow using Cadence tools to be taped out as an SoC with ARM-based NanoSoC platform.
@@
We evaluate cgra4ml on representative scientific workloads, including ResNet-50 for high-resolution image classification and PointNet for particle cloud inference. Results on AMD-Xilinx FPGAs and Cadence ASIC flows demonstrate that cgra4ml achieves competitive performance-per-watt, extensibility, and scalability, making it a robust and practical solution for real-time scientific inference at the edge.
2. Background and Related Work
This section outlines the requirements for a hardware-accelerated system development in scientific computing. We first introduce the need for neural networks in scientific computing and practical use cases. We then discuss two ways of implementing such networks: a layer-by-layer pipelined implementation where each layer becomes a pipeline stage in hardware and the accelerator-based implementation where all layers reuse the same reconfigurable hardware. hls4ml (FastML Team, 2024) and finn (Umuroglu et al., 2017) take the first approach, while several accelerators (Juracy et al., 2023) (Dhilleswararao et al., 2022) and frameworks such as Apache TVM with VTA accelerator (Moreau et al., 2018) take the second. We then outline the uniqueness and utility of a CGRA-type accelerator architecture to solve this problem and discuss related research in this area. Consequently, we summarize the needs of the scientific computing community that are not addressed by other generic frameworks as motivation for cgra4ml.
Figure 1. Data processing pipeline at Large Hadron Collider, CERN. The detectors at LHC produce data at a rate of 320 Tb/s during collisions. Due to the infeasibility of transferring data at such a high rate, the events are filtered through a multi-step pipeline. Small neural networks like autoencoders are implemented using hls4ml on radiation-hardened ASICs. More sophisticated models are implemented on FPGAs (L1) and servers (L2) to filter the data fully. With cgra4ml enabling larger models at low power, more processing can be moved to the edge (towards the left), resulting in more robust filtering.
Data processing pipeline at Large Hadron Collider, CERN.
2.1. Neural Networks for Scientific Computing
@@
Machine learning is fueling scientific discoveries in particle physics, materials, dark matter, cosmology, nuclear physics, biomedical engineering, and health monitoring (Deiana et al., 2022). Specific examples include bio-signal classification (Ahmad and Zafar, 2024), tracking magnetohydrodynamic (MHD) instabilities in fusion reactors (Wei et al., 2023), reinforcement learning for accelerator beam control (Duarte et al., 2022), pulse detection for anti-neutrino detection (Axani et al., 2024), and accelerator controls for beam loss deblending (Arnold et al., 2023). Scientific computing increasingly relies on neural networks to process data with requirements for extremely low latencies and high throughputs.
The Large Hadron Collider (LHC) has $\mathcal{O}(100M)$ individual sensors generating data for each of the 40 million proton beam collisions per second (Qu et al., 2022). LHC trigger systems filter events with sub-microsecond latency and 40 MHz throughput requirements using FPGAs (Duarte et al., 2018) as shown in Fig. 1. In each level, data is compressed, analyzed, and filtered. While traditional algorithms have been used for data filtering in the past, neural networks accelerated in FPGA and ASICs are replacing them. For example, neural network-based autoencoders are implemented as ASICs in the L0 trigger to compress and move data out of the sensors (Di Guglielmo et al., 2021). L1 uses more complex networks accelerated on FPGAs, and L2 triggers and offline reconstruction use even bigger models executed on GPUs. User-friendly frameworks for hardware implementation have fueled the wide adoption of small neural networks in lower-level triggers. hls4ml is one of the more popular among scientific edge accelerator frameworks.
Large models, such as Particle Transformer (Qu et al., 2022) for particle jet classification, are being explored by the physics community. However, for low-latency FPGA trigger use, they have not yet shown compelling gains, so practical deployments remain GPU-based. This is partly a chicken-and-egg issue. Since hls4ml implementation of models requires more and more resources with increasing model size, almost all the effort from the scientific computing community goes into making their models smaller to fit within an FPGA. Multi-FPGA routes, such as AIgean (Tarafdar et al., 2022), point to the possibility of using larger models but are comparatively cumbersome for trigger pipelines.
2.2. Dataflow-style NN Implementation: hls4ml, finn
@@
Figure 2. Layer-by-layer dataflow-style implementation of neural networks using hls4ml. hls4ml offers multiple backends, one per family of devices. Each backend is a collection of layers defined as templated HLS. Each neural network layer is implemented as a separate datapath. Xilinx DMA IPs are connected to the HLS4ML design to move data in and out. finn has a similar dataflow-style implementation. In contrast, cgra4ml reuses the same CGRA multiple times to process a layer.
Layer-by-layer spatial implementation of neural networks using hls4ml
hls4ml is a Python library developed low-latency, high-throughput edge machine learning inference (FastML Team, 2024). hls4ml primarily supports models built and trained using QKeras, a library from Google for low-precision machine learning (QKeras Team, 2024). hls4ml converts each layer of the model into a customized high-level synthesis (HLS) hardware block and connects them together to form an IP for the entire model (see Figure 2). The user controls reuse factor per layer, which defines the number of times a multiplier/accumulator is reused. The reuse factor provides a way to control a layer’s resource usage and performance. Balancing performance and resource utilization is highly dependent on the specific target FPGA and involves iterative manual tuning of reuse factors on a per-layer basis. This optimization effort is non-trivial and takes a significant time and effort.
Although the layer-by-layer approach may offer advantages in latency minimization, the architecture suffers from scaling issues as the size of its target NN increases. Because each layer is implemented with a unique datapath, larger NNs with tens to hundreds of layers quickly become too complex to implement on even the largest of hardware devices.
As the name implies, hls4ml requires a high-level synthesis (HLS) tool flow to generate the hardware accelerator. hls4ml has different HLS backends (AMD Xilinx FPGA, Altera FPGA, Siemens Catapult, etc.) with different levels of support. The most common target is AMD Xilinx Vitis/Vivado HLS, though other backends exist in various forms (Altera HLS Compiler, Siemens Catapult HLS, etc.). The dependence on HLS restricts the user to the supported vendors and requires HLS tool-specific reimplementation of each hardware layer, complicating the maintenance and verification. hls4ml’s support for ASIC implementation is primarily through Catapult HLS, which currently only supports a limited subset of layers.
A common hls4ml design flow requires scientists to generate an IP for the neural network, and then engineers write firmware, perform hardware-software integration, verification, and debugging, which is an onerous task. There are efforts to streamline this process for the scientists. For example, hls4ml provides a PYNQ overlay, a Python-based runtime for AMD Xilinx Zynq FPGAs. This workflow allows the physicists to focus on building and training low-precision models for their custom applications without spending months on hardware design, implementation, and verification. Yet, by and large, the ML IP core integration for production-ready designs remains a separate process from hls4ml.
finn (Umuroglu et al., 2017) is another popular open-source FPGA accelerator framework that uses a similar methodology. finn solely targets FPGAs though its general methodology could be ported for other hardware backends. finn shares many of the benefits and drawbacks of hls4ml, while being especially optimized for extremely low bitwidths, such as binary neural networks (Borras et al., 2022). The layer-by-layer approach is ideal for minimizing latency, since it eliminates the need to transfer partial outputs to/from the off-chip memory. It also makes the NN implementation extremely modular, allowing any layer to be added in any order. Yet, finn is primarily focused on generating an IP core that must be separately integrated into a larger system.
2.3. DNN Accelerators
Figure 3. Positioning CGRA4ML in the space of ML-to-FPGA frameworks from the perspective of the scientific computing community. hls4ml (FastML Team, 2024) is the most popular tool in the scientific computing community, as it is user-friendly and supports the dataflow-style implementation of models of varying bitwidths and layers to an extent. finn (Umuroglu et al., 2017) implements models in a similar dataflow fashion, but excels at very low bitwidths. Traditional AI accelerators are suitable for large models with 8+ bit quantization. cgra4ml aims to fill the gap by making models with sub 8-bit quantization that are too big to implement with hls4ml and finn.
Layer-by-layer spatial implementation of neural networks using hls4ml
Hundreds of accelerator architectures have been designed for deep neural network inference in the past decade (Peccerillo et al., 2022; Juracy et al., 2023; Dhilleswararao et al., 2022). This has been the default approach for the implementation of DNNs on hardware. These accelerators can be categorized by their data reuse pattern as weight, input, output and row stationary and by their runtime flexibility into systolic arrays, CGRAs, and microcode processors.
Caffeine (Zhang et al., 2019) offers an HLS systolic array to process CNNs from the Caffe ML framework using weight major and input major mappings. Eyeriss (Chen et al., 2017) is an energy-efficient ASIC designed to accelerate CNNs using a 2D array of 168 fairly complex processing elements, each with a 16-bit multiply-accumulate, scratchpad, and a controller. ShiDianNao (Du et al., 2015) uses a 2D mesh of functional units optimized towards the 2D feature maps of convolutional layers. Each processing element executes multiplications, additions, and comparisons using 16-bit fixed point arithmetic. The kernel elements are shifted right to left and up to down and accumulate locally.
While these architectures demonstrate strong performance and efficient data reuse, they often lack the broader system-level features required for scientific edge applications. Many of these accelerators are optimized for fixed model types, offering limited programmability and making it challenging to re-target them for newer or more diverse neural networks. Integration into embedded systems is frequently an afterthought, as they rarely include clean, general-purpose interfaces for control or data exchange with a microprocessor, which complicates deployment in real-time, heterogeneous environments. Additionally, verification support is typically constrained to smoke tests, presenting challenges for researchers and engineers who need robust testing and validation pipelines for use in high-stakes scientific contexts. Our framework addresses these gaps by combining performance-efficient hardware with runtime programmability, a microprocessor-friendly interface for seamless system integration, and a comprehensive verification suite tailored to the demanding requirements of scientific edge computing.
2.4. CGRA
@@
Coarse-Grained Reconfigurable Arrays (CGRA) is a type of architecture that has seen recent resurgence in industry and academic designs. Commercial CGRAs include Samsung Reconfigurable Processor (Kim et al., 2012) and the Renesas STP Reconfigurable Processor (Corporation, ). CGRA-ME (Chin et al., 2017) is an exemplary academic framework that supports a subset of CGRA designs.
CGRAs were originally envisioned as coarse-grained FPGAs whose programmable elements work at the word-level instead of FPGA bit-level programming (Liu et al., 2019). Early CGRAs were classified based on their integration into the processor. Tightly coupled CGRAs integrate into a processor data path and are executed as a custom instruction, e.g., Chess (Marshall et al., 1999), MATRIX (Mirsky and DeHon, 1996), and DySer (Govindaraju et al., 2012). Loosely coupled CGRAs act more as an accelerator that executes alongside the processor, executing in tandem with the processor and communicating via on-chip interconnect. Examples of loosely-coupled CGRAs include PipeRench (Goldstein et al., 2000), MorphoSys (Singh et al., 2000), CHARM (Cong et al., 2012), and FPCA (Cong et al., 2014). cgra4ml can be characterized as a loosely coupled CGRA as it implements sub-tasks of the neural network alongside a CPU with a focus on enabling hardware/software partitioning across the processor and processor.
Modern neural network accelerators often exhibit architectural similarities to CGRAs, particularly in their use of spatial dataflows, on-chip communication networks, and pipelined compute units. Like CGRAs, many NN accelerators rely on configurable interconnects and local memory structures to maximize data reuse and throughput. However, neural network accelerators are typically more narrowly optimized for specific computational patterns and may hardwire aspects of control or scheduling for performance. In contrast, CGRAs are generally more flexible—supporting a broader range of computational kernels and allowing dynamic reprogramming of data paths at runtime. With cgra4ml, we aim to navigate the design space between these two classes: delivering an architecture that captures the efficiency and specialization of NN accelerators while retaining the programmability and reuse benefits of CGRAs.
In the recent literature, OpenCGRA (Tan et al., 2020) introduced a unified, open-source stack to model, test, simulate, and characterize CGRAs, with LLVM compiler passes and PyMTL-based Verilog generation, such that designers can quickly explore heterogeneous tiles and full EDA flows in one place. AURORA (Tan et al., 2021) then automated CGRA co-design: starting from a generic architecture template, it jointly explores composition of functional units, memory/interconnect, and loop optimizations using an architecture-aware simulated annealing search. VecPAC (Tan et al., 2023) proposes a hybrid CGRA supporting different datatypes with scalar tiles plus configurable vector tiles. PICACHU (Qin et al., 2025) introduces a plug-in CGRA for LLM nonlinear ops (GELU/Softmax/LayerNorm, etc.) with domain-specialized PEs, precision-aware design, and seamless integration with a systolic array. ML-CGRA (Luo et al., 2025) is an MLIR-based end-to-end framework that keeps ML-level semantics and adds passes lower whole models onto CGRAs. AHA introduces an agile CGRA-compiler co-design flow where three domain specific languages: PEak for PEs, Lake for memories, and Canal for interconnects, auto-generate both RTL and compiler collateral from a single source of truth, so the compiler updates automatically as hardware evolves. While such CGRA frameworks primarily target architecture exploration and compiler lowering, cgra4ml’s use case is complementary and deployment-oriented. It delivers vendor-agnostic SystemVerilog RTL and a production C runtime verified in-the-loop via DPI-C, packaged as AXI IP with DMA that can be integrated with SoC platforms such as Zynq, NanoSoC and Ibex, enabling rapid, reproducible deployment in heterogeneous SoCs.
Table 1. Comparing cgra4ml with existing ML-to-FPGA/ASIC frameworks. As a deployment-oriented framework, cgra4ml’s unique feature is its comprehensive verification approach, which tests the user’s model and the production runtime on the generated hardware via DPI-C with randomized SoC-level congestion emulation. In addition, we provide a superset of features from other frameworks in one place for the scientific computing community
This work hls4ml (FastML Team, 2024) VTA (Moreau et al., 2018) FINN (Umuroglu et al., 2017) Vitis AI (AMD-Xilinx, 2024)
@@
DLP (Mathworks, 2024) OpenVino (Intel, 2024) Caffeine (Zhang et al., 2019) CGRA-ME (Chin et al., 2017) FlexCNN (Basalama et al., 2023)
Open-source
Sub-8-bit quantization support
Larger neural networks support
Xilinx FPGA support
Intel FPGA support
ASIC Implementation
Basic verification
Generate production-ready runtime
Verification with runtime
2.5. ML to FPGA/ASIC Frontend Frameworks